How many ounces of a 15% alcohol solution and a 42% alcohol solution must be contained to obtain 69 ounces of a 24% solution?
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

x=15%
69-x=42%
0.15x+(69-x)(0.42)=69(.24)=16.56
0.15x+28.98-0.42x=16.56
-0.27x=-12.42
x=46 ounces of 15%
69-x=23 ounces of 42%
===============
Way 2:
24% is 1/3 the way between 15 and 42%
Therefore, 2/3 is the 15% and 1/3 the 42%, and that is 46 and 23 ounces.

Answer by ikleyn(52788)

.
How many ounces of a 15% alcohol solution and a 42% alcohol solution must be contained to obtain 69 ounces of a 24% solution?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

For detailed solutions of the similar problems see the lessons

   Solving typical word problems on mixtures for solutions,  Problem 1,  and

   Typical word problems on mixtures from the archive,  Problems 2 and 3

in this site.